OpenStudy (anonymous):

If f(x)=x-x^3 and h doesn't equal 0, find the following: f(x+h)

OpenStudy (anonymous):

f(x) = x(1 - x^2) replace (x + h) wherever the function has x: f(x + h) = (x + h) [1 - (x + h)^2]